The active mass of 7.0 g nitrogen in a 2.0 l container would be
1
Expert's answer
2018-03-13T08:50:10-0400
Solution:
Moles of N2 = m/M = 7/28 = 0.25 mol
Active mass = concentratione in mol ( L-1)
Active mass = 0.25 mol / 2 L = 0.125 mol/L.
Answer: 0.125 mol/L.